1.     Find the price elasticity of supply at price of Birr 8 if the supply function is:

Q = 25 – 4P + P2

What is elasticity at Birr 4 and at Birr 5? Interpret the result.


1
Expert's answer
2021-11-30T10:10:05-0500

"Q=25-4P+P^2"

Elasticity of supply:

"=\frac{P}{Q(P)}\frac{dQ}{dP}"

"=\frac{P}{25-4P+P^2}(-4+2P)"

"=\frac{-4P+2P^2}{25-4P+P^2}"

At P= Birr 8,

"=\frac{-4(8)+2(8^2)}{25-4(8)+8^2}=1.88"

Price elasticity of supply being 1.88 implies that supply is elastic.


At P=Birr 4,

Elasticity of supply:

"=\frac{-4(4)+2(4^2)}{25-4(4)+4^2}=0.64"

Elasticity of supply being 0.64 implies that supply is inelastic.


At P=Birr 5,

Elasticity of supply:

"=\frac{-4(5)+2(5^2)}{25-4(5)+5^2}=1"

Elasticity of supply being 1 implies that supply is unit elastic.